OpenStudy (anonymous):

Solve this system of equations using substitution. a+b+c=180 b=3a c=5a

OpenStudy (anonymous):

a+3a+5a=180

OpenStudy (anonymous):

9a=180

OpenStudy (anonymous):

a=20

OpenStudy (anonymous):

so i basically just plug in the numbers

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

o ok thx u sooooo much

OpenStudy (anonymous):

now i understand lol

OpenStudy (anonymous):

yeah np smurf

OpenStudy (anonymous):

so would i divide for c or wut

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

itz askin for c now

OpenStudy (lalaly):

a=20 so b=3*20 = 60 c=5*20 = 100

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

so i multiply it ok

OpenStudy (karatechopper):

@mighty the answer is a=20

OpenStudy (anonymous):

yeah lana is right :)

OpenStudy (anonymous):

thank u

OpenStudy (karatechopper):

welcome:)

OpenStudy (lalaly):

:)